# Burst-tree structure and higher-order temporal correlations

### v2

@@ Abstract
Understanding characteristics of temporal correlations in time series is crucial for developing accurate models in natural and social sciences. The burst-tree decomposition method was recently introduced to reveal higher-order temporal correlations in time series in a form of an event sequence, in particular, the hierarchical structure of bursty trains of events for the entire range of timescales [Jo et al., Sci. Rep. 10, 12202 (2020)]. Such structure has been found to be simply characterized by the burst-merging kernel governing which bursts are merged together as the timescale for defining bursts increases. In this work, we study the effects of kernels on the higher-order temporal correlations in terms of burst size distributions, memory coefficients for bursts, and the autocorrelation function. We employ several kernels, including the constant, sum, product, and diagonal kernels as well as those inspired by empirical results. We generically find that kernels with preferential merging lead to heavy-tailed burst size distributions, while kernels with assortative merging lead to positive correlations between burst sizes. The decaying exponent of the autocorrelation function depends not only on the kernel but also on the power-law exponent of the interevent time distribution. In addition, thanks to the analogy to the coagulation process, analytical solutions of burst size distributions for some kernels could be obtained. Our findings may shed light on the role of burst-merging kernels as underlying mechanisms of higher-order temporal correlations in time series.
@@ I Introduction
Complex systems, which are ubiquitous in both natural and social systems, are distinguished by their intricate networks of interactions between individual components, leading to emergent complex dynamics [1, 2, 3, 4, 5, 6, 7]. These systems often exhibit bursty behaviors, where periods of intense activity are interspersed with periods of quiescence [8, 9]. Bursty dynamics have been observed in a wide range of systems, including physical [10, 11, 12, 13], biological [14, 15, 16], and social systems [17, 18, 19, 9, 20, 21]. Therefore, understanding temporal correlations in bursty time series of various complex systems is crucial for developing accurate models in natural and social sciences.
@@
Recently, Jo et al. introduced a novel method called burst-tree decomposition to analyze temporal correlations in the event sequence [20]. This method represents an event sequence in terms of the burst tree, which is decomposed into a set of IETs and an ordinal burst tree. The ordinal burst tree exactly captures the hierarchical structure of correlations between IETs in a systematic way. It turns out that several time series data from diverse backgrounds with different IET distributions show the similar burst-tree structure in terms of heavy-tailed burst size distributions and positive correlations between consecutive burst sizes for a wide range of timescales. The latter was quantified in terms of a memory coefficient for bursts. Originally, the memory coefficient has been proposed for detecting correlations between two consecutive IETs [17]. The burst-tree structure could be simply summarized by a burst-merging kernel governing which bursts are merged together as the timescale increases. The estimated burst-merging kernels from the data have revealed the preferential and assortative mixing structure of bursts. Then a model adopting a simple mathematical form of the burst-merging kernel based on the empirical analysis was studied to successfully generate temporal correlations in the original time series, providing a powerful tool for modeling and analyzing complex systems.
@@
The modeling approach using burst-merging kernels is analogous to the coagulation process by which unit particles are merged according to some selection rules or kernels, to become clusters [33, 34, 35, 36, 37, 38, 39, 40, 41]. Different approaches have been taken to derive cluster size distributions as well as power-law exponent values in the cases of power-law distributions of cluster sizes. In particular, Wattis employed a discrete deterministic mean-field approach to derive the analytical solution of the cluster size distribution for three solvable kernels, namely, constant, sum, and product kernels [41]. Similar results for the diagonal kernel have been obtained in Ref. [40]. Since unit particles and clusters in the coagulation process respectively correspond to individual events and bursts in the event sequence, the analytical solutions of cluster size distributions in the coagulation process can be translated into the burst size distributions of the event sequence generated using the same kernels. It implies that the analytical solutions of cluster size distributions in Refs. [41] and [40] can be used to confirm the burst size distributions obtained when the constant, sum, product, and diagonal kernels are employed as the burst-merging kernels.
@@
Despite the importance of burst-merging kernels in time series analysis, their impact on higher-order temporal correlations remains poorly understood. To address this issue, we conduct a systematic comparison of various kernels, including the constant, sum, product, and diagonal kernels as well as those inspired by empirical results. By analyzing the event sequences generated using these kernels and power-law IET distributions as inputs, we investigate burst size distributions, memory coefficients for bursts, and ACFs. We generically find that kernels with preferential merging lead to the heavy-tailed burst size distributions, while kernels with assortative merging lead to positive correlations between burst sizes. The decaying exponent of the autocorrelation function depends not only on the kernel but also on the power-law exponent of the IET distribution. These findings provide new insights into the role of burst-merging kernels as underlying mechanisms of higher-order temporal correlations in time series.
@@ III.1 Kernel-based modeling
As the empirical results indicate both preferential and assortative merging of bursts [20], we employ three more kernels dictating both preferential and assortative merging behaviors:
@@ III.3 Analytical results of burst size distributions
Figure 3: Simulation results of burst size distributions $Q_{\Delta t}(b)$, memory coefficients for bursts $M_{\Delta t}$ in Eq. (10), and autocorrelation functions $A(t_{\rm d})$ in Eq. (11) (from left to right) obtained from $50$ event sequences of $n=5\cdot 10^{5}$ events generated using $K_{\rm const}$, $K_{\rm sum}$, and $K_{\rm prod}$ [Eqs. (1)–(3)] (from top to bottom). In panels (a, d, g), we show $Q_{\Delta t}(b)$ for several values of $\Delta t$ in the case with the IET exponent $\alpha=2.5$, and $\langle b\rangle_{\Delta t}$ is the average burst size of $Q_{\Delta t}(b)$. Black lines with exponent values in panels (d, g) are for guiding eyes. Black lines in panels (c, f, i) show fitting ranges and results of the decay exponent $\gamma$ in Eq. (12), summarized in Table 1. Error bars in $M_{\Delta t}$ and $A(t_{\rm d})$ denote standard errors.
@@ IV Numerical results
Let us first look at the cases with constant, sum, and product kernels as shown in Fig. 3. The burst size distributions at several timescales, $Q_{\Delta t}(b)$, are exponential for $K_{\rm const}$ as expected. For $K_{\rm sum}$ and $K_{\rm prod}$, $Q_{\Delta t}(b)$s at several values of ${\Delta t}$ show power-law scaling regimes with exponents $3/2$ and $5/2$, respectively. Note that burst size distributions for $K_{\rm prod}$ contain small peaks at the very large $b$ of the order of $n$, which are not shown in Fig. 3(g). These very large bursts correspond to the gels in the coagulation process. As mentioned, all of constant, sum, and product kernels may not lead to positive correlations between consecutive burst sizes. Indeed, we observe $M_{\Delta t}\approx 0$ for these kernels. However, in the case of $K_{\rm const}$, we find $M_{\Delta t}\approx 0.2$ for small values of $\Delta t$ [Fig. 3(b)]. To explain this observation, we remind that for small $\Delta t$, there are a number of bursts of sizes $1$ or $2$. If these bursts were located in the sequence of burst sizes in a random manner, $M_{\Delta t}$ would be negligible. Instead, bursts of size $2$ seem to be located next to each other more often than expected, possibly due to the rule of randomly assigning left and right children nodes to the merged bursts. For example, if four bursts of sizes $1$, $1$, $2$, and $2$ are merged to two bursts of size $3$, which in turn are merged to a burst of size $6$, the sequence of burst sizes of $\{1,2,2,1\}$ may happen with a probability of $1/4$, eventually resulting in the positive $M_{\Delta t}$. This general argument however does not apply to the cases with $K_{\rm sum}$ and $K_{\rm prod}$ because the preferential merging forces consecutive burst sizes to differ from each other very quickly, leading to the negligible values of $M_{\Delta t}$. The negative values of $M_{\Delta t}$ for a very large $\Delta t$ are due to the finite-size effects as a few very big bursts and other relatively small bursts remain for such a large $\Delta t$. We finally remark that the IET exponent $\alpha$ affects only the timescales of the burst tree $\mathcal{T}$ not the ordinal burst tree $\mathcal{G}$; the smaller value of $\alpha$ implies a wider range of IETs and timescales. Therefore, the curve of $M_{\Delta t}$ for smaller $\alpha$ can be seen as a stretched version of that for larger $\alpha$.
@@
In Fig. 3(c, f, i), we find that the ACFs decay algebraically with the time lag $t_{\rm d}$. The estimated values of the decay exponent $\gamma$ are summarized in Table 1. The event sequences generated using $K_{\rm const}$ can be considered as results of the renewal process, in which there is no correlations between IETs. In that case, $\gamma$ has been known to be a function of $\alpha$ only; precisely, $\gamma=2-\alpha$ for $21<\alpha<$ 2 and $\gamma=\alpha-2$ for $\alpha>2$ [43]. However, numerical studies have shown strong finite-size effects on the mentioned scaling relation [42]. Considering such finite-size effects, the estimated values of $\gamma$ in Table 1 seem to be overall consistent with the analytical results of $\gamma$. In the case of $K_{\rm sum}$, $\beta=3/2$ implies long-range temporal correlations between consecutive IETs, leading to different values of $\gamma$ than those for $K_{\rm const}$. Although there are no analytical solutions of the ACF for $K_{\rm sum}$, one can refer to the recent analytical result of $\gamma$ as a function of $\alpha$ and $\beta$ in Ref. [30]; for the range of $21<\beta<$ 2, $\gamma$ is an increasing function of $\alpha$. Thus, it is qualitatively consistent with the numerical values of $\gamma$ for $K_{\rm sum}$ in Table 1. We remark that since the model studied in Ref. [30] considers only one timescale $\Delta t$, its analytical results might systematically deviate from numerical results of our model considering the entire range of $\Delta t$.
@@
The estimated value of $\beta$ for $K_{{\rm diag\text{-}pre}}$ is $\approx 0.85$, indicating a heavier tail than that for $K_{\rm diag}$. This can be understood by considering the fact that $K_{{\rm diag\text{-}pre}}$ has a preferential factor $b$ multiplied to $K_{\rm diag}$. However, such a numerical result is not consistent with the analytical solution in Eq. (16), requiring further investigation in the future. Next, by comparing $\beta\approx 1$ for $K_{{\rm sum\text{-}diag}}$ to $\beta=3/2$ for $K_{\rm sum}$, we observe that the tails of burst size distributions get heavier due to the assortative merging with $c_{1}>0$ in Eq. (6). We also find $\beta\approx 1.28$ for $K_{\rm emp}$. It turns out that estimated values of $\beta$ for these three kernels are smaller than $2$, allowing us to expect that $\gamma$ increases with $\alpha$ [30], which is indeed the case as shown in Table 1. The scaling behaviors for the case with $K_{\rm emp}$ appear most similar to those with $K_{\rm sum}$, except for the value of $\gamma$ when $\alpha=3.5$ is used.
@@ V Conclusion
Despite the usefulness of our framework using burst-merging kernels in generating or analyzing higher-order temporal correlations, underlying mechanisms behind such kernels are not yet fully understood. The kernels are used to determine which bursts are merged to become bigger bursts, rather than explain the reasons for such a determination at the fundamental level. Thus, it would be interesting if one could devise more mechanistic processes, e.g., based on queuing models [8, 45, 46] or self-exciting point processes [47, 48, 49, 44], and translate them into the burst-merging kernels. This topic is left for future works.

### v3

@@ Abstract
Understanding characteristics of temporal correlations in time series is crucial for developing accurate models in natural and social sciences. The burst-tree decomposition method was recently introduced to reveal temporal correlations in time series in a form of an event sequence, in particular, the hierarchical structure of bursty trains of events for the entire range of timescales [Jo et al., Sci. Rep. 10, 12202 (2020)]. Such structure cannot be solely captured by the interevent time distribution but can show higher-order correlations beyond interevent times. It has been found to be simply characterized by the burst-merging kernel governing which bursts are merged together as the timescale for defining bursts increases. In this work, we study the effects of kernels on the higher-order temporal correlations in terms of burst size distributions, memory coefficients for bursts, and the autocorrelation function. We employ several kernels, including the constant, sum, product, and diagonal kernels as well as those inspired by empirical results. We generically find that kernels with preferential merging lead to heavy-tailed burst size distributions, while kernels with assortative merging lead to positive correlations between burst sizes. The decaying exponent of the autocorrelation function depends not only on the kernel but also on the power-law exponent of the interevent time distribution. In addition, thanks to the analogy to the coagulation process, analytical solutions of burst size distributions for some kernels could be obtained. Our findings may shed light on the role of burst-merging kernels as underlying mechanisms of higher-order temporal correlations in time series.
@@ I Introduction
Complex systems, which are ubiquitous in both natural and social systems, are distinguished by their intricate networks of interactions between individual components, leading to emergent complex dynamics [1, 2, 3, 4, 5, 6, 7]. These systems often exhibit bursty behaviors, where periods of intense activity are interspersed with periods of quiescence [8, 9]. Bursty dynamics have been observed in a wide range of systems, including physical [10, 11, 12, 13], biological [14, 15, 16], and social systems [17, 18, 19, 20, 21, 22, 9, 23, 24, 25]. Therefore, understanding temporal correlations in bursty time series of various complex systems is crucial for developing accurate models in natural and social sciences.
@@
Recently, Jo et al. introduced a novel method called burst-tree decomposition to analyze temporal correlations in the event sequence [23]. This method represents an event sequence in terms of the burst tree, which is decomposed into a set of IETs and an ordinal burst tree. The ordinal burst tree exactly captures the hierarchical structure of correlations between IETs in a systematic way. It turns out that several time series data from diverse backgrounds with different IET distributions show the similar burst-tree structure in terms of heavy-tailed burst size distributions and positive correlations between consecutive burst sizes for a wide range of timescales. The latter was quantified in terms of a memory coefficient for bursts; note that the memory coefficient has been originally proposed for detecting correlations between two consecutive IETs [17]. Each node of the burst tree indicates a merge of two consecutive bursts that happens when the timescale $\Delta t$ exceeds the IET separating these two bursts. Which bursts are merged together as the timescale increases could be simply summarized by a function of burst sizes of two merged bursts, which is called a burst-merging kernel. Burst-merging kernels estimated from several empirical time series have revealed the preferential and assortative mixing structure of bursts in the data [23].
The burst-merging kernel can be interpreted as a selection rule of two bursts to be merged when generating the burst tree from the individual events. Once a burst-merging kernel is given either from the data or as a simple mathematical function, one can generate a burst tree, hence an event sequence. Indeed, a simplified version of empirically estimated kernels has been shown to successfully generate higher-order temporal correlations observed in the original time series [23]. Yet, analytical derivation of higher-order temporal correlations from the given kernel has been overwhelmingly difficult. Therefore, we take an alternative approach by studying the simplest, analytically tractable kernels first and then by considering the more complicated, realistic forms of kernels focused on preferential and assortative merging of bursts. By doing so, one can deepen the understanding of the effects of different features of the kernels on the higher-order temporal correlations in the generated time series. We remark that application of our method to the empirical data is beyond the scope of our current work.
@@
The modeling approach using burst-merging kernels is analogous to the coagulation process by which unit particles are merged according to some selection rules or kernels, to become clusters [37, 38, 39, 40, 41, 42, 43, 44, 45]. Different approaches have been taken to derive cluster size distributions as well as power-law exponent values in the cases of power-law distributions of cluster sizes. In particular, a discrete deterministic mean-field approach was employed to derive the analytical solution of the cluster size distribution for three solvable kernels, namely, constant, sum, and product kernels [45]. Similar results for the diagonal kernel have been obtained in Ref. [44]. Since unit particles and clusters in the coagulation process respectively correspond to individual events and bursts in the event sequence, the analytical solutions of cluster size distributions in the coagulation process can be translated into the burst size distributions of the event sequence generated using the same kernels. It implies that the analytical solutions of cluster size distributions in Refs. [45] and [44] can be used to confirm the burst size distributions obtained when the constant, sum, product, and diagonal kernels are employed as the burst-merging kernels.
@@
Despite the importance of burst-merging kernels in time series analysis, their impact on higher-order temporal correlations beyond the IET distribution remains poorly understood. To address this issue, we conduct a systematic comparison of various kernels, including the constant, sum, product, and diagonal kernels as well as those inspired by empirical results. By analyzing the event sequences generated using these kernels and power-law IET distributions as inputs, we investigate burst size distributions, memory coefficients for bursts, and ACFs. We generically find that kernels with preferential merging lead to the heavy-tailed burst size distributions, while kernels with assortative merging lead to positive correlations between burst sizes. The decaying exponent of the autocorrelation function depends not only on the kernel but also on the power-law exponent of the IET distribution. These findings provide new insights into the role of burst-merging kernels as underlying mechanisms of higher-order temporal correlations in empirical time series.
@@ III.1 Kernel-based modeling
The previous empirical results, e.g., using a Wikipedia editor’s edit sequence, indicate both preferential and assortative merging of bursts, namely, the empirically estimated kernels $K(b,b^{\prime})$ tend to increase with $b$ and $b^{\prime}$ and to show higher values around the diagonal line of $b=b^{\prime}$ than the off-diagonal regions [23]. Inspired by such results, we here employ three more kernels dictating both preferential and assortative merging behaviors:
@@ III.3 Analytical results of burst size distributions
Figure 3: Simulation results of burst size distributions $Q_{\Delta t}(b)$, memory coefficients for bursts $M_{\Delta t}$ in Eq. (10), and autocorrelation functions $A(t_{\rm d})$ in Eq. (11) (from left to right) obtained from $50$ event sequences of $n=5\cdot 10^{5}$ events generated using $K_{\rm const}$, $K_{\rm sum}$, and $K_{\rm prod}$ [Eqs. (1)–(3)] (from top to bottom). In panels (a, d, g), we show $Q_{\Delta t}(b)$ for several values of $\Delta t$ in the case with the IET exponent $\alpha=2.5$, and $\langle b\rangle_{\Delta t}$ is the average burst size of $Q_{\Delta t}(b)$. Black lines without or with exponent values in panels (a, d, g) are for guiding eyes. Black lines in panels (c, f, i) show fitting ranges and results of the decay exponent $\gamma$ in Eq. (12), summarized in Table 1. Error bars in $M_{\Delta t}$ and $A(t_{\rm d})$ denote standard errors.
@@ IV Numerical results
Let us first look at the cases with constant, sum, and product kernels as shown in Fig. 3. The burst size distributions at several timescales, $Q_{\Delta t}(b)$, are exponential for $K_{\rm const}$ as expected. For $K_{\rm sum}$ and $K_{\rm prod}$, $Q_{\Delta t}(b)$s at several values of ${\Delta t}$ show power-law scaling regimes with exponents $3/2$ and $5/2$, respectively. Note that for the case with $K_{\rm prod}$, since almost all events make very large bursts as partly shown in Fig. 2, the rest of events may not be enough to fully develop the power-law tails in the burst size distributions. Such large bursts correspond to the gels in the coagulation process and they appear in the burst size distributions as small peaks at the very large $b$ of the order of $n$, which are not shown in Fig. 3(g).
As mentioned, all of constant, sum, and product kernels may not lead to positive correlations between consecutive burst sizes. Indeed, we observe $M_{\Delta t}\approx 0$ for these kernels. However, in the case of $K_{\rm const}$, we find $M_{\Delta t}\approx 0.2$ for small values of $\Delta t$ [Fig. 3(b)]. To explain this observation, we remind that for small $\Delta t$, there are a number of bursts of sizes $1$ or $2$. If these bursts were located in the sequence of burst sizes in a random manner, $M_{\Delta t}$ would be negligible. Instead, bursts of size $2$ seem to be located next to each other more often than expected, possibly due to the rule of randomly assigning left and right children nodes to the merged bursts. For example, if four bursts of sizes $1$, $1$, $2$, and $2$ are merged to two bursts of size $3$, which in turn are merged to a burst of size $6$, the sequence of burst sizes of $\{1,2,2,1\}$ may happen with a probability of $1/4$, eventually resulting in the positive $M_{\Delta t}$. This general argument however does not apply to the cases with $K_{\rm sum}$ and $K_{\rm prod}$ because the preferential merging forces consecutive burst sizes to differ from each other very quickly, leading to the negligible values of $M_{\Delta t}$. The negative values of $M_{\Delta t}$ for a very large $\Delta t$ are due to the finite-size effects as a few very big bursts and other relatively small bursts remain for such a large $\Delta t$. We finally remark that the IET exponent $\alpha$ affects only the timescales of the burst tree $\mathcal{T}$ not the ordinal burst tree $\mathcal{G}$; the smaller value of $\alpha$ implies a wider range of IETs and timescales. Therefore, the curve of $M_{\Delta t}$ for smaller $\alpha$ can be seen as a stretched version of that for larger $\alpha$.
@@
In Fig. 3(c, f, i), we find that the ACFs decay algebraically with the time lag $t_{\rm d}$. We estimate the value of the decay exponent $\gamma$ by the ordinary least squares linear regression using the equation of $\ln A(t_{\rm d})=a_{0}-\gamma\ln t_{\rm d}$ with a constant $a_{0}$ for the scaling regime of $t_{\rm d}$. The results are summarized in Table 1 with the fitting ragne of $t_{\rm d}$. The event sequences generated using $K_{\rm const}$ can be considered as results of the renewal process, in which there is no correlations between IETs. In that case, $\gamma$ has been known to be a function of $\alpha$ only; precisely, $\gamma=2-\alpha$ for $21<\alpha<$ 2 and $\gamma=\alpha-2$ for $\alpha>2$ [47]. However, numerical studies have shown strong finite-size effects on the mentioned scaling relation [46]. Considering such finite-size effects, the estimated values of $\gamma$ in Table 1 seem to be overall consistent with the analytical results of $\gamma$. In the case of $K_{\rm sum}$, $\beta=3/2$ implies long-range temporal correlations between consecutive IETs, leading to different values of $\gamma$ than those for $K_{\rm const}$. Although there are no analytical solutions of the ACF for $K_{\rm sum}$, one can refer to the recent analytical result of $\gamma$ as a function of $\alpha$ and $\beta$ in Ref. [34]; for the range of $21<\beta<$ 2, $\gamma$ is an increasing function of $\alpha$. Thus, it is qualitatively consistent with the numerical values of $\gamma$ for $K_{\rm sum}$ in Table 1. We remark that since the model studied in Ref. [34] considers only one timescale $\Delta t$, its analytical results might systematically deviate from numerical results of our model considering the entire range of $\Delta t$.
@@
The estimated value of $\beta$ for $K_{{\rm diag\text{-}pre}}$ is $\approx 0.85$, indicating a heavier tail than that for $K_{\rm diag}$. This can be understood by considering the fact that $K_{{\rm diag\text{-}pre}}$ has a preferential factor $b$ multiplied to $K_{\rm diag}$. However, such a numerical result is not consistent with the analytical solution in Eq. (16), possibly due to the fact that the analytical solution was obtained for continuous variables of masses in the coagulation process in contrast to the discrete nature of the burst-merging process. This requires further investigation in the future. Next, by comparing $\beta\approx 1$ for $K_{{\rm sum\text{-}diag}}$ to $\beta=3/2$ for $K_{\rm sum}$, we observe that the tails of burst size distributions get heavier due to the assortative merging with $c_{1}>0$ in Eq. (6). We also find $\beta\approx 1.28$ for $K_{\rm emp}$. It turns out that estimated values of $\beta$ for these three kernels are smaller than $2$, allowing us to expect that $\gamma$ increases with $\alpha$ [34], which is indeed the case as shown in Table 1. The scaling behaviors for the case with $K_{\rm emp}$ appear most similar to those with $K_{\rm sum}$, except for the value of $\gamma$ when $\alpha=3.5$ is used.
@@ V Conclusion
Although we have focused on the preferential and assortative mixing structure of bursts, it is also possible that other kinds of features in burst-merging kernels may be observed in other empirical time series. Then our understanding of the simple kernels may serve as a reference to understand such novel features and their impact on the higher-order temporal correlations in such real-world phenomena.
Finally, despite the usefulness of our framework using burst-merging kernels in generating or analyzing higher-order temporal correlations, underlying mechanisms behind such kernels are not yet fully understood. The kernels are used to determine which bursts are merged to become bigger bursts, rather than explain the reasons for such a determination at the most fundamental level. Thus, it would be interesting if one could devise more mechanistic processes, e.g., based on queuing models [8, 49, 50] or self-exciting point processes [51, 52, 53, 48], and translate them into the burst-merging kernels. By doing so, one may relate elements in the mechanistic processes to the features of corresponding kernels.